OpenStudy (anonymous):

I'm trying to compute growth time.The problem says to use the natural logarithm on both sides of the equation "and use the property 1n M^p = p 1n M" How does that work? The equation is 1n 1.2 = n 1n 1.0075 Please help! Thx :)

OpenStudy (dumbcow):

is this the equation: \[\ln (1.2) =n*\ln (1.0075) \]

OpenStudy (anonymous):

Yes, then the next step shows \[n=\ln 1.2\div \ln 1.0075\] I'm trying to apply this equation in the book to a homework question. Thanks for helping out!

OpenStudy (dumbcow):

don;t know what you really need help with

OpenStudy (dumbcow):

the general exponential growth function is : \[P(t) = P_{0}*b^{kt}\]

OpenStudy (anonymous):

I don't know what that means. Thanks anyway I'm trying to understand what "use the property 1n M^p = p 1n M" when using the base e logarithm